CS 591 Randomized Algorithms, HW1

Prof. Jared Saia, University of New Mexico Due: September 14th, 2004

1. Exercise 1.2

2. Exercise 1.4

3. Exercise 1.9

4. Problem 1.1 part a (part b is extra credit)

5. Problem 1.8

6. Problem 1.9

7. Problem 1.10

8. Extra Credit Problem 1.6

9. Extra Credit: Problem 1.7